OpenStudy (idani):

write the expression as a single logarithm 7logby+6logbx

satellite73 (satellite73):

\[\huge 7\log_b(y)+6\log_b(x)\]?

satellite73 (satellite73):

\[n\log(x)=\log(x^m)\] is a start so for example \[7\log(y)=\log(y^7)\]

satellite73 (satellite73):

do the same for \[6\log(x)\]

satellite73 (satellite73):

let me know what you get

OpenStudy (idani):

Log(x^6)

satellite73 (satellite73):

yes ok one more step

satellite73 (satellite73):

\[\log(A)+\log(B)=\log(AB)\]

satellite73 (satellite73):

you have \[\large \log(y^7)+\log(x^6)\] combine them as above

OpenStudy (idani):

Log(x^6y^7)

satellite73 (satellite73):

bingo `

OpenStudy (idani):

Thank you so much ☺️☺️

satellite73 (satellite73):

yw

OpenStudy (silverspeedfighter):

Thats kinda Cool How you do that?

satellite73 (satellite73):

i have an advanced degree in elementary logariddims

OpenStudy (silverspeedfighter):

ok? @satellite73

satellite73 (satellite73):

also java http://openstudy.com/study#/updates/577c308de4b01e8189f355ce